OpenStudy (anonymous):

Bet no one can solve this one: Find the inverse laplace transform of log ((S^2+1)/(S^2))

hartnn (hartnn):

its easy

OpenStudy (anonymous):

Can you show me then? o-O

hartnn (hartnn):

do u know \(L[any function]= -1/t L[derivative of given function]\) ??

hartnn (hartnn):

sorry, those are inverse Laplace.

hartnn (hartnn):

\(L^{-1}[F(s)]= -1/t \quad L^{-1}[d/dt {F(s)}]\) did u know this ?

hartnn (hartnn):

\(L^{-1}[log ((S^2+1)/(S^2))]= -1/t \quad L^{-1}[d/dt[ {log ((S^2+1)/(S^2))]}]\)

hartnn (hartnn):

that should be d/ds so what is d/ds [log ((S^2+1)/(S^2))]

OpenStudy (anonymous):

Ahh I see. Thank you very much. I can do the rest from here. :)

hartnn (hartnn):

also verify your question...is it correct ? ot is there any square root sign ?

hartnn (hartnn):

hmm.. ok welcome ^_^